OpenStudy (anonymous):

whts 4 1/2 + 5 2/3=??

OpenStudy (anonymous):

4 1/2 + 5 2/3 = 4 3/6 + 5 4/6 = 9 7/6 = 10 1/6 --> ans:)

OpenStudy (anonymous):

Change the mixed fractions into improper fractions. To do this multiply the number in front by the denominator, add the numerator, and then put that number over the denominator. 9/2 + 17/3 9(3)/2(3) + 17(2)/13(2) 27/6 + 34/6 61/6 is your answer as an improper fraction, 10 1/6 as a mixed fraction.